# Modelling and Classifying the Components of a Literature Review

### v1

@@ E.5 Encoder-Decoder
For this category, we employed different versions of the T5 [139], including T5-base (222 million parameters), T5-large (770 million parameters), T5-xl (3 billion parameters), and T5-xxl (11 billion parameters). These models are built on a transformer architecture, wherein the encoder processes the input text, and the decoder generates the output text. T5 has been pretrained on the C4 corpus, a large dataset of text and code, using both supervised and self-supervised training methods. It has a context window of 512 tokens. Our decision to use T5 was driven by its superior performance compared to other encoder-decoder models [140, 141, 142].
@@ E.6 Encoder
In this category, all models, with the exception of BigBird [143], exhibit identical features, including a number of parameters of 110 million, a context length of 512 tokens, and the utilisation of a full attention mechanism. In contrast, BigBird distinguishes itself with 125 million parameters, an extended context length of 4,096 tokens, and the use of a sparse attention mechanism.
BERT base [144], hereafter referred to as BERT, was trained on a dataset containing 3.3 billion tokens sourced from Wikipedia and the Google Books Corpus. SciBERT case (SciBERT[84] was developed using 1.14 million scientific articles from Semantic Scholar282828 Semantic Scholar - (https://www.semanticscholar.org/), covering the biomedical and computer science domains, with a total of 3.1 billion tokens. BioBERT [145] was trained on an extensive collection of biomedical literature, including publications from PubMed292929 PubMed - (https://pubmed.ncbi.nlm.nih.gov/) and PMC303030 PMC - (https://pmc.ncbi.nlm.nih.gov/), and employs WordPiece [146] tokenization to efficiently manage a large vocabulary. On the other hand, BigBird was trained on a dataset of 1.5 billion tokens drawn from the Books Corpus and Wikipedia.

### v2

@@ E.5 Encoder-Decoder
For this category, we employed different versions of the T5 (Raffel et al., , 2020), including T5-base (222 million parameters), T5-large (770 million parameters), T5-xl (3 billion parameters), and T5-xxl (11 billion parameters). These models are built on a transformer architecture, wherein the encoder processes the input text, and the decoder generates the output text. T5 has been pretrained on the C4 corpus, a large dataset of text and code, using both supervised and self-supervised training methods. It has a context window of 512 tokens. Our decision to use T5 was driven by its superior performance compared to other encoder-decoder models (Garg et al., , 2021; Sarrouti et al., , 2022; Kementchedjhieva and Chalkidis, , 2023).
@@ E.6 Encoder
In this category, all models, with the exception of BigBird (Zaheer et al., , 2020), exhibit identical features, including a number of parameters of 110 million, a context length of 512 tokens, and the utilisation of a full attention mechanism. In contrast, BigBird distinguishes itself with 125 million parameters, an extended context length of 4,096 tokens, and the use of a sparse attention mechanism.
BERT base (Devlin, , 2018), hereafter referred to as BERT, was trained on a dataset containing 3.3 billion tokens sourced from Wikipedia and the Google Books Corpus. SciBERT case (SciBERT) (Beltagy et al., , 2019) was developed using 1.14 million scientific articles from Semantic Scholar [^28] Semantic Scholar - (https://www.semanticscholar.org/), covering the biomedical and computer science domains, with a total of 3.1 billion tokens. BioBERT (Lee et al., , 2020) was trained on an extensive collection of biomedical literature, including publications from PubMed [^29] PubMed - (https://pubmed.ncbi.nlm.nih.gov/) and PMC [^30] PMC - (https://pmc.ncbi.nlm.nih.gov/), and employs WordPiece (Wu, , 2016) tokenization to efficiently manage a large vocabulary. On the other hand, BigBird was trained on a dataset of 1.5 billion tokens drawn from the Books Corpus and Wikipedia.